OpenStudy (chrisga):

The table below shows some values of f(x) and g(x) for different values of x: x f(x) = 9x + 7 g(x) = 5x −2 −11 −1 −2 0 1 1 5 2 Complete the chart and determine the solution of the equation f(x) = g(x). x = −1 x = 0 x = 2 x = 25

OpenStudy (chrisga):

I already solved the graph x f(x) = 9x + 7 g(x) = 5x -2 -11 0.04 -1 -2 0.2 0 7 1 1 16 5 2 25 25 But im not sure is the answer C or D

OpenStudy (chrisga):

@LynFran

OpenStudy (lynfran):

ur g(x) is incorrect g(x)=5x when x=-2 , 5(-2)=-10 when x=-1 , 5(-1)= -5 when x=0 , 5(0)=0 when x=1 5(1)=5 when x=2 5(2)=10

OpenStudy (lynfran):

f(x)=g(x) 9x+7=5x 9x-5x=-7 4x=-7 x=-7/4... im not getting those answers are you sure those are the correct function for f(x)=9x+7 g(x)=5x ???

OpenStudy (chrisga):

yes

OpenStudy (chrisga):

x = −1 x = 0 x = 2 x = 25

OpenStudy (chrisga):

i might just put D